OpenStudy (anonymous):

Point G is the centroid of the triangle. The length of BG =10. What is the length of BG?

OpenStudy (anonymous):

30 15 20 18

OpenStudy (anonymous):

huh?? read the ques carefully! :D

OpenStudy (anonymous):

"The length of BG =10"

OpenStudy (anonymous):

yes

OpenStudy (rsadhvika):

you want to find BE ?

OpenStudy (anonymous):

yes want the know the length of BE

OpenStudy (rsadhvika):

hint : centroid divided the median in 2:1 ratio

OpenStudy (anonymous):

centroid divided the median in ration 2:1 so.. if bg=10 be= \[(2x \div3)\times10\]

OpenStudy (jiteshmeghwal9):

First of all u should know that the centroid of a triangle divides each median into ratio 2:1

OpenStudy (anonymous):

yes I know that

OpenStudy (anonymous):

There will be a ratio ratio 2:1 when cetroid divided the median.

OpenStudy (jiteshmeghwal9):

BG=2x=10 remaining part =x

OpenStudy (anonymous):

so do I do 10 times 2 which equals 20?

OpenStudy (anonymous):

no.. 10/2 =5

OpenStudy (anonymous):

No you will have BG+ GE= BE BG=10 GE=10/2 BE=? 10+10/2=BE 10+5=BE

OpenStudy (anonymous):

okay so the answer would be 15

OpenStudy (anonymous):

yup good job

OpenStudy (anonymous):

Thank you so much you guys it really helped a lot